OpenStudy (anonymous):

can someone show me hot to find the quotient for 8a^6/2a^3

jimthompson5910 (jim_thompson5910):

8 divided by 2 = 4

jimthompson5910 (jim_thompson5910):

a^6 divided by a^3 is a^3 (you subtract the exponents)

jimthompson5910 (jim_thompson5910):

so the final quotient is 4a^3

OpenStudy (anonymous):

oh awessome! i did it right!!!!

jimthompson5910 (jim_thompson5910):

ok that's great

OpenStudy (anonymous):

what about one like... -9m^7n^4/-3m^4n^3

jimthompson5910 (jim_thompson5910):

-9 divided by -3 = ???

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

m^7 divided by m^4 = ??

OpenStudy (anonymous):

i thought you subtract the exponents

jimthompson5910 (jim_thompson5910):

you do

jimthompson5910 (jim_thompson5910):

so what do you get

OpenStudy (anonymous):

3m^3n?????

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

YAYYYY! okay wait what about doing one likeee... simplifying (4p^3/3q^2)^2

jimthompson5910 (jim_thompson5910):

square each term

jimthompson5910 (jim_thompson5910):

so square 4p^3 to get ???

OpenStudy (anonymous):

16p^6

jimthompson5910 (jim_thompson5910):

now square 3q^2

OpenStudy (anonymous):

9q^4

OpenStudy (anonymous):

then 16 times 9?

jimthompson5910 (jim_thompson5910):

no it's a fraction remember?

OpenStudy (anonymous):

OHHHH

OpenStudy (anonymous):

waqit so just leave i as a fraction.. like 16p^6/9q^4

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

hmmmm okay. what about one that as simple as (-10)^3 just evaluating that..

OpenStudy (anonymous):

pr 2^-4

OpenStudy (anonymous):

or*

jimthompson5910 (jim_thompson5910):

(-10)^3 is just -1000

jimthompson5910 (jim_thompson5910):

you cube 10 and make it negative

OpenStudy (anonymous):

oh alright...so for the second one how would i do that

jimthompson5910 (jim_thompson5910):

for the second one you would use the idea that x^(-k) = 1/(x^k)

jimthompson5910 (jim_thompson5910):

so 2^(-4) = 1/(2^4) = ???

OpenStudy (anonymous):

wait..... what?

jimthompson5910 (jim_thompson5910):

In other words, I did this